OpenStudy (anonymous):

NEED HELP: The dimensions of a rectangle are 10 feet by 15 feet respectively. If each of the dimensions is increased by the same amount, the perimeter of the new rectangle is 28 feet more than the perimeter of the old rectangle. Find the amount increase.

OpenStudy (kropot72):

Original perimeter = 2 * 10 + 2* 15 = 50 ft Let x = amount increase New perimeter = 50 + 28 = 2(10 + x) + 2(15 + x) 78 = 50 + 4x 4x = 28 Can you find x now?

OpenStudy (anonymous):

x =7

OpenStudy (anonymous):

is that the final answer?

OpenStudy (kropot72):

The amount increase is 7 feet on each of the four sides, giving a total increase of 28 feet.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (kropot72):

You're welcome :)